OpenStudy (anonymous):

please help!! fan and medal!! b^-3=1/216 A. –6 B. 6 C. 36 D. 72

jagr2713 (jagr2713):

\[\huge \color {Blue}{Welcome \to OpenStudy!} \Huge \color{blue}{\star^{ \star^{\star:)}}}\Huge \color{red}{\star^{ \star^{\star:)}}}\]

OpenStudy (ribhu):

b^-1/3 = 1/b^3

OpenStudy (alexandervonhumboldt2):

b^-3=1/216 b^3=216 b=cube rootout of 216

OpenStudy (ribhu):

@AlexandervonHumboldt2 correct explanation.